# An Entropy-based Text Watermarking Detection Method

### v3

@@ 7 Conclusion
In this work, we proposed a text watermark detection algorithm called EWD based on token entropy. We gave theoretical analysis on the detection properties of our method in low-entropy scenarios with comparison of previous works. We also validated the detection performance of our method in the low-entropy code generation task with a number of weight functions. Furthermore, our method would achieve similar performance in high-entropy scenarios as the previous works. Experiments also showed that compared to previous methods, our method would not bring extra computational cost.
@@
The limitations of our method mainly include two parts. The first part is that the low-entropy datasets tested are limited. We only conducted experiments on two code generation datasets, and we did not conduct experiments on more low-entropy datasets. At the same time, our method should be theoretically effective for various types of watermark methods, such as token sampling-based methods. This paper did not try more types of methods, which is an important future work direction.
@@ Appendix A Proof of Type-II Error for EWD
Here we demonstrate how we obtained the theoretical Type-II error for EWD. First of all, we attempted to calculate the theoretical $\mu$ and $\sigma^{2}$ based on Eq.6. The theoretical $\mu$ is the weighted sum of green tokens in the watermarked text, using ${|s|_{G}}_{k}$ to represent the weighted number of green tokens from each token $k$,thus we can obtain:
$\mu$ $=\sum_{k}{|s|_{G}}_{k}$ (7)
$\mu$ $=\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ (8)
$=T\mathbb{E}\frac{\gamma\alpha SE}{1+(\alpha-1)\gamma}(SE-C_{0})$ (9)
@@
Given the simulated power law of spike entropy, we can obtain $E[SE^{2}]$ and $E[SE]$ as 0.377 and 0.608 respectively, producing $\mu$ equal to 5.79. And for $\sigma^{2}$, we first obtain the variance for each token $k$:
${\sigma_{k}}^{2}$ $=\mathbb{E}[({|s|_{G}}_{k}-\mathbb{E}[{|s|_{G}}_{k}])^{2}]$ (11)
$=\mathbb{E}[{|s|_{G}}_{k}^{2}]-\mathbb{E}[{|s|_{G}}_{k}]^{2}$ (12)
$=\mathbb{P}[k\in G]\cdot W(k)^{2}-\mathbb{P}[k\in G]^{2}\cdot W(k)^{2}$ (13)
@@
$\sigma^{2}$ $=\sum_{k}W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G])$ (15)
$=T\mathbb{E}[W(k)^{2}C_{1}SE(1-C_{1}SE)]$ (16)
$=TC_{1}\mathbb{E}[(SE-C_{0})^{2}SE(1-C_{1}SE)]$ (17)
@@
Given the simulated power law of spike entropy, we can obtain $E[SE^{4}]$ and $E[SE^{3}]$ as 0.24 and 0.158 respectively, producing $\sigma$ equal to 0.56. When the detection threshold $z$ is set as 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55), the probability of $|s|_{G}\sim\mathcal{N}(5.79,0.56)$ falling below 5.55 is 33.3%, $i.e.$ Type-II error.

### v4

@@ 7 Conclusion
In this work, we propose a text watermark detection method called EWD that fully considers the influence of token entropy. We first establish the positive influence-entropy relation in the detection process, then utilize entropy for weight customization to fully reflect the proposed relation. We provide theoretical analysis on the detection properties of different methods in low-entropy scenarios. Experiments validate the detection performance of our method in the low-entropy code generation task with a number of weight functions. Furthermore, in high-entropy scenarios, our method would achieve similar performance as previous works in detection accuracy, detection speed and robustness.
@@
Our method mainly includes two limitations. The first one is that the low-entropy datasets tested are limited. We utilize two code generation datasets for evaluation. In the future, we aim to include more low-entropy tasks and datasets.
Secondly, our method should be theoretically effective for various types of watermark methods, such as token sampling-based methods. In the future, we aim to implement our method on more types of text watermarking frameworks.
Acknowledgement
We sincerely appreciate the insightful comments from all reviewers. Their help has greatly contributed to the improvement of this work.
The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 14222922, RGC GRF 2151185).
@@ Appendix A Proof of Type-II Error for EWD
Here we demonstrate how we obtain the theoretical Type-II error for EWD. First of all, we attempt to calculate the theoretical $\mu$ and $\sigma^{2}$ based on Eq. (6). The theoretical $\mu$ is the weighted sum of green tokens in the watermarked text, using ${|s|_{G}}_{k}$ to represent the weighted number of green tokens from each token $k$,thus we can obtain:
$\mu$ $=\sum_{k}{|s|_{G}}_{k},$ (7)
$=\sum_{k}\mathbb{P}[k\in G]\cdot W(k),$ (8)
$=T\mathbb{E}\frac{\gamma\alpha SE}{1+(\alpha-1)\gamma}(SE-C_{0}),$ (9)
@@
Given the simulated power law of spike entropy, we can obtain $E[SE^{2}]$ and $E[SE]$ as 0.377 and 0.608, respectively, producing $\mu$ equal to 5.79. And for $\sigma^{2}$, we first obtain the variance for each token $k$:
${\sigma_{k}}^{2}$ $=\mathbb{E}[({|s|_{G}}_{k}-\mathbb{E}[{|s|_{G}}_{k}])^{2}],$ (11)
$=\mathbb{E}[{|s|_{G}}_{k}^{2}]-\mathbb{E}[{|s|_{G}}_{k}]^{2},$ (12)
$=\mathbb{P}[k\in G]\cdot W(k)^{2}-\mathbb{P}[k\in G]^{2}\cdot W(k)^{2},$ (13)
@@
$\sigma^{2}$ $=\sum_{k}W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G]),$ (15)
$=T\mathbb{E}[W(k)^{2}C_{1}SE(1-C_{1}SE)],$ (16)
$=TC_{1}\mathbb{E}[(SE-C_{0})^{2}SE(1-C_{1}SE)],$ (17)
@@
Given the simulated power law of spike entropy, we can obtain $E[SE^{4}]$ and $E[SE^{3}]$ as 0.24 and 0.158, respectively, producing $\sigma$ equal to 0.56. When the detection threshold $z^{\prime}$ is set as 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55), the probability of $|s|_{G}\sim\mathcal{N}(5.79,0.31)$ falling below 5.55 is 33.4%, $i.e.$, the Type-II error.